OpenStudy (anonymous):

how do I solve: . A 5-kg fish swimming at 1 m/s swallows an absentminded 1-kg fish swimming toward it at a velocity that brings both fish to a halt. Show that the speed of the smaller fish before lunch was 5 m/s.

OpenStudy (abhisar):

Hello @beccamarx19 !

OpenStudy (abhisar):

Let the speed of the small fish before lunch be x m/s

OpenStudy (abhisar):

Now by using Principle of conservation of linear momentum (m1v1 = m2v2)

OpenStudy (abhisar):

-5 * 1 + x * 1 = 0

OpenStudy (abhisar):

Solving this we get x = 5

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

i got that never mind

OpenStudy (abhisar):

are uh sure ?

OpenStudy (anonymous):

if it is -5*1 it equals -5 then x*1 =0 -5+5 equal 0

OpenStudy (abhisar):

No, -5*1 + 1*x = 0, then, -5=-x, then, 5=x

OpenStudy (anonymous):

meaning x has to be 5 in order for it all to equal 0

OpenStudy (abhisar):

yes

OpenStudy (anonymous):

right i think i got it thank you

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.

OpenStudy (abhisar):

Feel free to tag me in if u hav more doubts \(\color{green}{\huge\ddot\smile}\)

OpenStudy (anonymous):

how do i do that i posted another question already

OpenStudy (abhisar):

just add @ with anyones name u want to tag like if u want to tag me write @abhisar